# Faster Convergence of Riemannian Stochastic Gradient Descent with Increasing Batch Size

### v3

@@ Appendix B Objective Function Value against SFO Complexity
Figure 6: Objective function value against SFO complexity on COIL100 (PCA), MNIST (PCA), MovieLens-1M (LRMC), and Jester (LRMC) in order from left to right.
Appendix C Proof of The Propositions
C.1 Proof of Lemma 1
@@
We define $\mathcal{F}_{t}\coloneq\sigma[x_{0},(\xi_{j,0})_{j},(\xi_{j,1})_{j},\ldots,(\xi_{j,t-1})_{j}]$, where $x_{0}$ is the initial point of RSGD. For measurable function $f$, $\mathbb{E}[f|\mathcal{F}_{t}]$ denotes the conditional expectation of $\mathcal{F}_{t}$.
Proof.
Considering Assumption 1, we start with
@@
$\mathbb{E}[\langle\mathrm{grad}f(x_{t}),\mathrm{grad}f_{B_{t}}(x_{t})\rangle_{x_{t}}|\mathcal{F}_{t+1}]=\langle\mathrm{grad}f(x_{t}),\mathbb{E}[\mathrm{grad}f_{B_{t}}(x_{t})|\mathcal{F}_{t+1}]\rangle_{x_{t}}=\|\mathrm{grad}f(x_{t})\|_{x_{t}}^{2}\qquad\text{a.s.}$
Taking the total expectation for the above equations, we have
@@
C.2 Proof of Lemma 2
Proof.
@@ C.2 Proof of Lemma 2
By taking the summation on both sides of the inequality for Lemma 1 and evaluating it using the above inequality, we obtain
@@
C.3 Proof of Theorem 1
Proof.
We set $b_{t}=b$ in Lemma 2.
[Constant LR equation 1]
From Lemma 2, we have
@@ C.3 Proof of Theorem 1
[Diminishing LR equation 2]
From Lemma 2, we have
@@
[Cosine Annealing LR equation 3]
From Lemma 2, we have
@@
[Polynomial Decay LR equation 4]
We start by recalling the Riemann integral (e.g., Rudin (1976)) of $g(t)\coloneq(1-t)^{p}\geq 0(0\leq t\leq 1)$. $U(P_{T})\coloneq\sum_{t=0}^{T-1}\frac{1}{T}(1-\frac{t}{T})^{p}$ is an upper Riemann sum with the function $g$ and a partition $P_{T}\coloneq\{0,\frac{1}{T},\cdots,\frac{T-1}{T},1\}$ due to $g$ being monotone decreasing. From the definition of an Riemann integral (e.g., Rudin (1976, Section 6)), if $\int_{0}^{1}g(t)dt$ exists, then
@@
Similarly, if we define $L(P_{T})$ as $\sum_{t=1}^{T}\frac{1}{T}\left(1-\frac{t}{T}\right)^{p}$, then $L(P_{T})$ is a lower Riemann sum with the function $g$ and a partition $P_{T}$, and the supremum of $L(P_{T})$ equals $\int_{0}^{1}g(t)dt$. Considering
@@
Therefore, considering Lemma 2,
@@
$=\eta_{\max}^{2}-\eta_{\min}^{2}+\left(\eta_{\min}+\frac{2\eta_{\min}(\eta_{\max}-\eta_{\min})}{p+1}+\frac{(\eta_{\max}-\eta_{\min})^{2}}{2p+1}\right)T,$
we have
@@
C.4 Proof of Theorem 2
Proof.
The evaluations in all these cases are based on Lemma 2 and use estimations of $\sum_{t=0}^{T-1}\eta_{t}$ in the proof of Theorem 1.
[Exponential Growth BS equation 5 and Constant LR equation 1]
@@ C.4 Proof of Theorem 2
[Exponential Growth BS equation 5 and Diminishing LR equation 2]
Using equation 9, we have
@@
[Exponential Growth BS equation 5 and Cosine Annealing LR equation 3]
From equation 9 and $|\cos x|\leq 1$, we have
@@
[Exponential Growth BS equation 5 and Polynomial Decay LR equation 4]
From equation 9 and $t\leq T$, we have
@@
[Polynomial Growth BS equation 6 and Constant LR equation 1]
@@
where, for $c>1$, Riemann zeta function $\zeta(c)<\infty$; $\zeta(c)$ is monotone decreasing on $c\in(1,\infty)$ and $\underset{c\to\infty}{\lim}\zeta(c)=1$.
[Polynomial Growth BS equation 6 and Diminishing LR equation 2]
From equation C.4, we have
@@
[Polynomial Growth BS equation 6 and Cosine Annealing LR equation 3]
From equation C.4 and the previous analysis for Cosine Annealing LR equation 3, we have
@@
[Polynomial Growth BS equation 6 and Polynomial Decay LR equation 4]
From equation C.4 and the previous analysis for Polynomial Decay LR equation 4, we have
@@
C.5 Proof of Theorem 3
Proof.
The evaluations in all these cases are based on Lemma 2. We start with an exponential growth BS and a warm-up LR. From $l,l_{w}\in\mathbb{N}$ and $l_{w}\geq l$, there exist $\alpha,\beta\in\mathbb{N}\cup\{0\}$ such that $l_{w}=\alpha l+\beta$. Note that we define summations from 0 to 1-1 as 0. Furthermore, from $\forall u,v>0:0\leq u^{2}+v^{2}+uv$ holds.
[LR Decay Part: Constant equation 1]
@@ C.5 Proof of Theorem 3
[LR Decay Part: Diminishing equation 2]
@@
[LR Decay Part: Cosine Annealing equation 3]
@@
[LR Decay Part: Polynomial Decay equation 4]
As with the proof of polynomial decay LR equation 4 in Theorem 1, we have
@@
On the other hand, we consider a polynomial growth BS and a warm-up LR.
[LR Decay Part: Constant equation 1]
@@
[LR Decay Part: Diminishing equation 2]
@@
[LR Decay Part: Cosine Annealing equation 3]
@@
[LR Decay Part: Polynomial Decay equation 4]
Considering the proof of Polynomial Decay LR equation 4 in Theorem 1, we have
@@
C.6 Proof of Theorem 4
Proof.
[LR Decay Part: Constant equation 1]
Considering C.5, we have
@@ C.6 Proof of Theorem 4
[LR Decay Part: Diminishing equation 2]
@@
[LR Decay Part: Cosine Annealing equation 3]
Doing the same with C.5, we have
@@
[LR Decay Part: Polynomial Decay equation 4]
@@
Appendix D Objective Function Values in Section 4.1 and Section 4.2
@@
Figure 7: Objective function value against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 in order from left to right on COIL100 dataset.
@@
Figure 8: Objective function value against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 in order from left to right on MNIST dataset.
@@
Figure 9: Objective function value against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 in order from left to right on MovieLens-1M dataset.
@@
Figure 10: Objective function value against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 in order from left to right on Jester dataset.
The performance in terms of the objective function value against the number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 on the COIL100, MNIST, MovieLens-1M, and Jester datasets are shown in Figures 7, 8, 9, and 10, respectively.
Appendix E Additional Numerical Results

### v4

@@ A.2 Objective Function Value versus SFO Complexity
Figure 7: Objective function value (loss) versus SFO complexity on COIL100 (PCA), MNIST (PCA), MovieLens-1M (LRMC), and Jester (LRMC) datasets in order from left to right.
For the COIL100 (PCA) and MovieLens-1M (LRMC), performance was better with an increasing BS than with a constant BS. Although the differences in the objective function values for the MNIST (PCA) and Jester (LRMC) are small (one possible reason for this is that the objective function may be flat around the optimal solution), the performance with an increasing BS was equal to or better than that with a constant BS. A more detailed discussion of this hypothesis is presented in Section 4.3.
Appendix B Trade-offs for Each Hyperparameter of Exponential Growth BS Scheduler
We calculate the trade-offs for each hyperparameter of the exponential growth BS.
$\gamma$]: From Theorem 3.3, $\|G_{T}\|^{2}=O\left(1+\frac{\gamma}{\gamma-1}\right)$ and $\text{SFO}_{\epsilon}^{\text{incr}}=O(\gamma^{M})$ hold, and $\gamma^{M}$ is clearly monotonically increasing with respect to $\gamma$. Let $f(\gamma)\coloneq 1+\frac{\gamma}{\gamma-1}$. Because $^{\prime}(\gamma)=\frac{-1}{(\gamma-1)^{2}}<0, ff$ is monotonically decreasing. Thus, for a smaller gradient norm, $\gamma$ should be set to a larger value, while for a smaller SFO complexity, $\gamma$ should be set to a smaller value.
$b_{0}$]: From Theorem 3.3, $\|G_{T}\|^{2}=O\left(1+b_{0}^{-1}\right)$ and $\text{SFO}_{\epsilon}^{\text{incr}}=O\left(\frac{b_{0}^{3}}{b_{0}^{2}-1}\right)$ hold. $1+b_{0}^{-1}$ is clearly monotonically decreasing. Let $g(x)\coloneq\frac{x^{3}}{x^{2}-1}\ (x>1)$. Because $^{\prime}(x)=\frac{x^{2}(x-\sqrt{3})(x+\sqrt{3})}{(x^{2}-1)^{2}}, gg$ is monotonically decreasing for $1<x\leq\sqrt{3}$, and monotonically increasing for $x\geq\sqrt{3}$. Now, because $x>1$ and $b_{0}\in\mathbb{N}$, it suffices to consider only the case $x\geq 2$, in which $g$ is monotonically increasing. Summarizing the above and considering $b_{0}\in\text{dom}\ g$, we find that, for a smaller gradient norm, $b_{0}$ should be set to a larger value, while for a smaller SFO complexity, $b_{0}$ should be set to a smaller value.
[$M$]: Recall $T=MK$. From Theorem 3.3, $\|G_{T}\|^{2}=O\left(\frac{1}{T}+\frac{K}{T}\right)=O\left(\frac{1}{MK}+\frac{K}{MK}\right)=O\left(M^{-1}\right)$, and $\text{SFO}_{\epsilon}^{\text{incr}}=O\left(\frac{\gamma^{M}}{M}\right)$ hold. $M^{-1}$ is clearly monotonically decreasing. Let $h(x)\coloneq\frac{\gamma^{x}}{x}\ (x\geq 1)$. Because $^{\prime}(x)=\frac{\gamma^{x}(x\ln\gamma-1)}{x^{2}}, hh$ is monotonically increasing for $x\geq\frac{1}{\ln\gamma}$ when $\gamma\geq e$. Thus, when we set $\gamma\geq e$ ($\gamma$ was set to 3 in our experiments except for the far right plot in Figure 6), we find that, for a smaller gradient norm, $M$ should be set to a larger value, while for a smaller SFO complexity, $M$ should be set to a smaller value.
Finally, we provide a remark on the setting of $b_{0}$ in our experiments. Although the initial BS ($b_{0}$) appears different for each dataset, the values were similarly determined using the same method to enable a comparison of the effectiveness between a small constant BS, a large constant BS, and an increasing BS. The $N$ was set to the total number of data points in each dataset. For a large initial BS, it was set to $3^{k}$, where $k$ is the largest integer such that $3^{k}<N$. For a medium-size initial BS, it was set to $3^{k-3}$. For the smallest initial BS, it was set to $3^{k-5}$. For example, because $N=7200$ for the COIL100 dataset, the large, the medium-size small, and the smallest initial BSs were set to $3^{8}=6561,3^{5}=243,3^{3}=27$, respectively.
Appendix C Our Proposed Optimization Problem
Our proposed optimization problem:
minimize $f(w)\coloneq\frac{1}{N}\sum_{j=1}^{N}\sqrt{|\langle x_{j},w\rangle|},$
subject to $w\in\text{S}^{n-1}\coloneq\{w\in\mathbb{R}^{n}\mid\|w\|=1\}.$
$\{x_{j}\}_{j=1}^{N}$ is a dataset uniformly sampled from $\text{S}^{n-1}$. $N$ is a total number of data points in the dataset, and $n$ is a dimension of each one. The objective function defined above has unbounded gradient norms, which allows us to experimentally verify that our theoretical results are indeed applicable to functions with unbounded gradient norms. In our experiment, $R_{x}(d)\coloneq\frac{x+d}{\|x+d\|}$ was used as a retraction for $\text{S}^{n-1}$.
Proposition C.1.
The gradient norm of the objective function for this problem $f$ is not bounded.
Proof C.2.
When $\langle x_{j},w\rangle\not=0\ (^{\forall}j\in\{1,\ldots,N\})$,
$\nabla f(w)=\frac{1}{N}\sum_{j=1}^{N}\text{Sign}(\langle x_{j},w\rangle)|\langle x_{j},w\rangle|^{-\frac{1}{2}}x_{j}$
is obtained by
$\frac{\partial f}{\partial w_{i}}(w)$ $=\frac{1}{N}\sum_{j=1}^{N}\frac{\partial}{\partial w_{i}}\sqrt{|\langle x_{j},w\rangle|}=\frac{1}{N}\sum_{j=1}^{N}|\langle x_{j},w\rangle|^{-\frac{1}{2}}\text{Sign}(\langle x_{j},w\rangle)\frac{\partial}{\partial w_{i}}\sum_{k=1}^{N}x_{j}^{k}w_{k}$
$=\frac{1}{N}\sum_{j=1}^{N}\text{Sign}(\langle x_{j},w\rangle)|\langle x_{j},w\rangle|^{-\frac{1}{2}}x_{j}^{i},$
where $w=(w_{1},\cdots,w_{n})^{\top},x_{j}=(x_{j}^{1},\cdots,x_{j}^{n})^{\top}\in\text{S}^{n-1}\subset\mathbb{R}^{n}$. Thus, the gradient on $\text{S}^{n-1}$ is
$\text{grad}f(w)=(I-ww^{\top})\nabla f(w)=\frac{1}{N}\sum_{j=1}^{N}\text{Sign}(\langle x_{j},w\rangle)|\langle x_{j},w\rangle|^{-\frac{1}{2}}(I-ww^{\top})x_{j}.$
Fix one data $x_{i}$ such that $x_{i}\not=0$. Let $u\in\text{S}^{n-1}$ be a vector obtained by orthonormalizing $x_{i}$ via the Gram-Schmidt process. We consider a sequence $(w_{t})_{t\in\mathbb{N}}$ such that $w_{t}\coloneq\frac{x_{i}}{t\|x_{i}\|}+\sqrt{1-\frac{1}{t^{2}}}u$, which lies in $\text{S}^{n-1}$ because
$\|w_{t}\|^{2}=\frac{1}{t^{2}\|x_{i}\|^{2}}\|x_{i}\|^{2}+\left(1-\frac{1}{t^{2}}\right)\|u\|^{2}+\frac{2}{t\|x_{i}\|}\sqrt{1-\frac{1}{t^{2}}}\langle x_{i},w\rangle=1.$
By simple calculations,
$w_{t}w_{t}^{\top}$ $=\left(\frac{x_{i}}{t\|x_{i}\|}+\sqrt{1-\frac{1}{t^{2}}}u\right)\left(\frac{x_{i}^{\top}}{t\|x_{i}\|}+\sqrt{1-\frac{1}{t^{2}}}u^{\top}\right)$
$=\frac{x_{i}x_{i}^{\top}}{t^{2}\|x_{i}\|^{2}}+\frac{1}{t^{2}\|x_{i}\|^{2}}\sqrt{1-\frac{1}{t^{2}}}(x_{i}u^{\top}+ux_{i}^{\top})+\left(1-\frac{1}{t^{2}}\right)uu^{\top}$
$\overset{t\to\infty}{\longrightarrow}uu^{\top}$
and
$\langle x_{j},w_{t}\rangle=\left\langle x_{j},\frac{x_{i}}{t\|x_{i}\|}+\sqrt{1-\frac{1}{t^{2}}}u\right\rangle=\frac{\langle x_{j},x_{i}\rangle}{t\|x_{i}\|}+\sqrt{1-\frac{1}{t^{2}}}\langle x_{j},u\rangle$
hold. Now, we define a set $\Lambda$ satisfying $\{x_{j}\}_{j\in\Lambda}\subset\{x_{j}\}_{j=1}^{N}$ such that $\langle x_{j},u\rangle=0$. We find $\langle x_{j},w_{t}\rangle=\frac{\langle x_{j},x_{i}\rangle}{t\|x_{i}\|}(^{\forall}j\in\Lambda)$, and $\Lambda\not$$\varnothing$ because of $i\in\Lambda$. Note that, for all $j\not\in\Lambda$, $\langle x_{j},u\rangle\not=0$ holds, which yields $\langle x_{j},w_{t}\rangle\not=0$. From these observations, if $\langle x_{j},x_{i}\rangle\not=0\ (^{\forall}j\in\Lambda)$,
$\|\text{grad}f(w_{t})\|$ $=\left\|\frac{1}{N}\sum_{j=1}^{N}\text{Sign}(\langle x_{j},w_{t}\rangle)|\langle x_{j},w_{t}\rangle|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|$
$=\frac{1}{N}\left\|\sum_{j=1}^{N}|\langle x_{j},w_{t}\rangle|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|$
$\geq\frac{1}{N}\left\|\sum_{j\in\Lambda}|\langle x_{j},w_{t}\rangle|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|-\frac{1}{N}\left\|\sum_{j\not\in\Lambda}|\langle x_{j},w_{t}\rangle|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|$
$=\frac{1}{N}\sqrt{\frac{t\|x_{i}\|}{\langle x_{j},x_{i}\rangle}}\left\|\sum_{j\in\Lambda}(I-w_{t}w_{t}^{\top})x_{j}\right\|-\frac{1}{N}\left\|\sum_{j\not\in\Lambda}\left|\frac{\langle x_{j},x_{i}\rangle}{t\|x_{i}\|}+\sqrt{1-\frac{1}{t^{2}}}\langle x_{j},u\rangle\right|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|$
$\geq\frac{1}{N}\sqrt{\frac{t\|x_{i}\|}{\langle x_{j},x_{i}\rangle}}\left\|\sum_{j\in\Lambda}(I-w_{t}w_{t}^{\top})x_{j}\right\|-\frac{1}{N}\sum_{j\not\in\Lambda}\left|\frac{\langle x_{j},x_{i}\rangle}{t\|x_{i}\|}+\sqrt{1-\frac{1}{t^{2}}}\langle x_{j},u\rangle\right|^{-\frac{1}{2}}\left\|(I-w_{t}w_{t}^{\top})x_{j}\right\|$
$\overset{t\to\infty}{\longrightarrow}\left(\infty-\frac{1}{N}\sum_{j\not\in\Lambda}\left|\langle x_{j},u\rangle\right|^{-\frac{1}{2}}\left\|(I-uu^{\top})x_{j}\right\|=\right)\infty$
holds; even if $\langle x_{j},x_{i}\rangle=0\ (^{\exists}j\in\Lambda)$, for all $t\in\mathbb{N}$,
$\|\text{grad}f(w_{t})\|$ $\geq\frac{1}{N}\left\|\sum_{j\in\Lambda}|\langle x_{j},w_{t}\rangle|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|-\frac{1}{N}\left\|\sum_{j\not\in\Lambda}|\langle x_{j},w_{t}\rangle|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|$
$=\left(\infty-\frac{1}{N}\left\|\sum_{j\not\in\Lambda}|\langle x_{j},w_{t}\rangle|^{-\frac{1}{2}}(I-w_{t}w_{t}^{\top})x_{j}\right\|=\right)\infty$
holds. These complete the proof.
C.1 Gradient Norm versus SFO Complexity for Our Proposed] Optimization Problem
Figure 8: Norm of objective function gradient versus SFO complexity. Plot on left (resp. right) shows norm of objective function gradient for our proposed problem when $(N,n)=(7200,1024)$ (resp. when $(N,n)=(60000,1024)$).
As shown in the plot of the norm of objective function gradient versus SFO complexity 8 for our proposed problem, performance was better with an increasing BS than with either a small or large constant BS. This result can be explained by Theorem 4.1, which shows that an increasing BS reduces SFO complexity compared with a constant BS.
Appendix D Proof of our Lemma and Theorems
@@
We define $\mathcal{F}_{t}\coloneq\sigma[x_{0},(\xi_{j,0})_{j},(\xi_{j,1})_{j},\ldots,(\xi_{j,t-1})_{j}]$, where $x_{0}$ is the initial point of RSGD. For a measurable function $f$, $\mathbb{E}[f|\mathcal{F}_{t}]$ denotes the conditional expectation with respect to $\mathcal{F}_{t}$. Note that, from Assumption 2, $\mathbb{E}_{\xi\sim\Pi}\left[\|\mathrm{grad}f_{\xi}(x)-\mathrm{grad}f(x)\|_{x}^{2}\right]=\mathbb{V}_{\xi\sim\Pi}(\mathrm{grad}f_{\xi}(x))\leq\sigma^{2}$ holds. The following result serves as a preliminary for Lemma 3.1.
Lemma D.1 (Descent Lemma).
Let $(x_{t})_{t}$ be a sequence generated by RSGD and $(\eta_{t})_{t}$ be a positive-valued sequence. Then, under Assumptions 2.1 and 2, we obtain
Proof D.2.
Under Assumption 2.1, we start with
@@
$\mathbb{E}[\langle\mathrm{grad}f(x_{t}),\mathrm{grad}f_{B_{t}}(x_{t})\rangle_{x_{t}}|\mathcal{F}_{t+1}]=\langle\mathrm{grad}f(x_{t}),\mathbb{E}[\mathrm{grad}f_{B_{t}}(x_{t})|\mathcal{F}_{t+1}]\rangle_{x_{t}}=\|\mathrm{grad}f(x_{t})\|_{x_{t}}^{2}\quad\text{a.s.}$
Taking the total expectation of the above equations, we have
@@
D.1 Proof of Lemma 3.1
On the basis of Lemma D.1, Lemma 3.1 is proven as follows.
Proof D.3.
@@
By taking the summation on both sides of the inequality for Lemma D.1 and evaluating it using the above inequality, we obtain
@@
D.2 Proof of Theorem 3.2
Proof D.4.
We set $b_{t}=b$ in Lemma 3.1.
[Constant LR (1)]
From Lemma 3.1, we have
@@
[Diminishing LR (2)]
From Lemma 3.1, we have
@@
[Cosine Annealing LR (3)]
From Lemma 3.1, we have
@@
[Polynomial Decay LR (4)]
We start by recalling the Riemann integral of $g(t)\coloneq(1-t)^{p}\geq 0$ $(0\leq t\leq 1)$. $U(P_{T})\coloneq\sum_{t=0}^{T-1}\frac{1}{T}(1-\frac{t}{T})^{p}$ is an upper Riemann sum for the function $g$ over the partition $P_{T}\coloneq\{0,\frac{1}{T},\cdots,\frac{T-1}{T},1\}$ due to $g$ being monotonically decreasing. From the definition of a Riemann integral, if $\int_{0}^{1}g(t)dt$ exists, then
@@
Similarly, if we define $L(P_{T})$ as $\sum_{t=1}^{T}\frac{1}{T}\left(1-\frac{t}{T}\right)^{p}$, then $L(P_{T})$ is a lower Riemann sum for the function $g$ over the partition $P_{T}$, and the supremum of $L(P_{T})$ equals $\int_{0}^{1}g(t)dt$. Considering
@@
Therefore, considering Lemma 3.1,
@@
$=\eta_{\max}^{2}-\eta_{\min}^{2}+\left(\eta_{\min}+\frac{2\eta_{\min}(\eta_{\max}-\eta_{\min})}{p+1}+\frac{(\eta_{\max}-\eta_{\min})^{2}}{2p+1}\right)T$
hold, which yields
@@
D.3 Proof of Theorem 3.3
Proof D.5.
The evaluations in all the following cases are based on Lemma 3.1 and use estimations of $\sum_{t=0}^{T-1}\eta_{t}$ in the proof of Theorem 3.2.
[Exponential Growth BS (5) and Constant LR (1)]
@@
[Exponential Growth BS (5) and Diminishing LR (2)]
Using (9), we have
@@
[Exponential Growth BS (5) and Cosine Annealing LR (3)]
From (9) and $|\cos x|\leq 1$, we have
@@
[Exponential Growth BS (5) and Polynomial Decay LR (4)]
From (9) and $t\leq T$, we have
@@
[Polynomial Growth BS (6) and Constant LR (1)]
@@
where, for $c>1$, Riemann zeta function $\zeta(c)<\infty$; $\zeta(c)$ is monotonically decreasing on $c\in(1,\infty)$ and $\underset{c\to\infty}{\lim}\zeta(c)=1$.
[Polynomial Growth BS (6) and Diminishing LR (2)]
From (D.5), we have
@@
[Polynomial Growth BS (6) and Cosine Annealing LR (3)]
From (D.5) and the previous analysis for Cosine Annealing LR (3), we have
@@
[Polynomial Growth BS (6) and Polynomial Decay LR (4)]
From (D.5) and the previous analysis for Polynomial Decay LR (4), we have
@@
D.4 Proof of Theorem 3.4
Theorem D.6 (Detailed Version of Theorem 3.4).
We consider BSs (5) and (6) and warm-up LRs (7) and (8) with decay parts given by (1), (2), (3), or (4) under the assumptions of Lemma 3.1. Then, the following holds for both constant and increasing BSs.
  • Decay part: Diminishing (2)
  • Decay part: Otherwise (1), (3), (4)
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants that do not depend on $T$.
Proof D.7.
The evaluations in all these cases are based on Lemma 3.1. We start with an exponential growth BS and a warm-up LR. From $l,l_{w}\in\mathbb{N}$, and $l_{w}\geq l$, there exist $\alpha,\beta\in\mathbb{N}\cup\{0\}$ such that $l_{w}=\alpha l+\beta$. Note that we define summations from 0 to 1-1 as 0. Furthermore, from $\forall u,v>0:0\leq u^{2}+v^{2}+uv$ holds.
[LR Decay Part: Constant (1)]
@@
[LR Decay Part: Diminishing (2)]
@@
[LR Decay Part: Cosine Annealing (3)]
@@
[LR Decay Part: Polynomial Decay (4)]
As with the proof for a polynomial decay LR (4) in Theorem 3.2, we have
@@
Next we consider a polynomial growth BS and a warm-up LR.
[LR Decay Part: Constant (1)]
@@
[LR Decay Part: Diminishing (2)]
@@
[LR Decay Part: Cosine Annealing (3)]
@@
[LR Decay Part: Polynomial Decay (4)]
Considering the proof of the polynomial decay LR (4) in Theorem 3.2, we have
@@
D.5 Proof of Theorem 3.5
Theorem D.8 (Detailed Version of Theorem 3.5).
We consider a constant BS ($b_{t}=b>0$) and warm-up LRs (7) and (8) with decay parts given by (1), (2), (3), or (4) under the assumptions of Lemma 3.1. Then, we obtain
  • Decay part: Diminishing (2)
  • Decay part: Otherwise (1), (3), (4)
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants that do not depend on $T$.
Proof D.9.
[LR Decay Part: Constant (1)]
Considering D.4, we have
@@
[LR Decay Part: Diminishing (2)]
@@
[LR Decay Part: Cosine Annealing (3)]
Doing the same with D.4, we have
@@
[LR Decay Part: Polynomial Decay (4)]
@@
Appendix E Objective Function Values in Sections 4.1 and 4.2
@@
Figure 9: Objective function value (loss) versus number of iterations for LRs (1), (2), (3), and (4) in order from left to right on COIL100 dataset (PCA).
@@
Figure 10: Objective function value (loss) versus number of iterations for LRs (1), (2), (3), and (4) in order from left to right on MNIST dataset (PCA).
@@
Figure 11: Objective function value (loss) versus number of iterations for LRs (1), (2), (3), and (4) in order from left to right on MovieLens-1M dataset (LRMC).
@@
Figure 12: Objective function value (loss) versus number of iterations for LRs (1), (2), (3), and (4) in order from left to right on Jester dataset (LRMC).
The performance in terms of the objective function value versus the number of iterations for LRs (1), (2), (3), and (4) on the COIL100, MNIST, MovieLens-1M, and Jester datasets are shown in Figures 9, 10, 11, and 12, respectively. Although the differences in the objective function values are small (one possible reason for this is that the objective function may be flat around the optimal solution), the performance with an increasing BS was equal to or better than that with a constant BS. A more detailed discussion of this hypothesis is provided in Section 4.3.
Appendix F Convergence Criteria in Table 1
In previous studies and in our work, three types of convergence criteria were used:
  1. 1.
    the objective function value $\mathbb{E}[f(x_{T})-f^{\star}]$,
  2. 2.
    the gradient value $\underset{t\to\infty}{\lim}\mathbb{E}[\mathrm{grad}f(x_{t})]=0$ or $\|G_{T}\|^{2}$,
  3. 3.
    (extended) Riemannian distance between $x_{T}$ and an optimal solution $x^{\star}$: $\mathbb{E}[\|\Delta_{T}\|]\coloneq\mathbb{E}[\|R_{x^{\star}}^{-1}(x_{T})\|]$.
Criterion (3) applies because $\mathbb{E}[\|\Delta_{T}\|]=\mathbb{E}[d(x_{T},x^{\star})]$ holds when $R=\text{Exp}$.
Appendix G Additional Numerical Results